[Gautam, Johannes, Koji, Annalisa]
Tonight we increased the power of the PSL laser and we achieved the lock of both arms with high power.
The AUX beam alignment to the Y arm was recovered and the PLL restored (using the Marconi as LO).
We made a quick measurement of the phase noise and the results will be posted tomorrow.
The beam on the PSL has been blocked, as well as the AUX beam on the AS table. The Marconi has been switched off.
gautam:
- Before turning up PSL power, I placed a block in front of MC refl to avoid any PD burning. Replaced HR Y1 2" optic with the usual 10% reflective BS to direct MC REFL to the locking PD.
- Waveplate was rotated back to 180 deg (original position before the vent). After optimizing PMC transmission, I measured 1.05 W going into the IMC (pre-vent value was 1.07 W, prolly within power meter absolute accuracy).
- IMC autolocker restored to usual high power version on megatron.
- There seems to be some kind of vacuum interlock in effect that prevents me from opening the PSL shutter via EPICS - I had to toggle the position on the shutter controller under the table. After tonight's work, I returned the controller to the NC state, to avoid any further interference with this interlock code that may prevent pumping in the AM.
- PLL gain was re-adjusted to achieve maximum stability (judged by eye) of the beat-note in lock triggered on the Marconi LO signal. Alignment onto the NF beatPD was also tweaked to squeeze out as much beat as possible.
- The main objective tonight was to send AUX beam in, recover transmission beat, scan the AUX frequency, and resolve some peaks (MAX HOLD scanning technique, magnitude only for now, no phase info). Thanks to JE's expert fiber alignment and beatnote maximization, we achieved this
. Annalisa will post a plot tmr.
- For unknown reasons, the Y arm ASS does not maximize TRY. So we are in the unfortunate situation of neither arm having a working ASS servo. To be worked on later.
|